OpenStudy (anonymous):

What is the value of the expression –5x^4 when x = –2? -40 -80 40 80

OpenStudy (lucaz):

plug in the value of x in the expression and calculate

OpenStudy (lucaz):

-5(-2)^4

OpenStudy (lucaz):

(-2)^4 = (-2)^2*(-2)^2 = 4*4 = 16 -5(16) = -80